Three different types of candy that cost $1.29, $1.79, and
$2.39 per pound are to be mixed to produce 14 pounds of
candy worth $23.86. If there is to be twice as much of the
$1.29 candy as the $2.39 candy, how much of each type
should be mixed together?

Answer by lwsshak3(11628)

Three different types of candy that cost $1.29, $1.79, and
$2.39 per pound are to be mixed to produce 14 pounds of
candy worth $23.86. If there is to be twice as much of the
$1.29 candy as the $2.39 candy, how much of each type
should be mixed together?
**
let x=pounds of 2.39 candy
2x=pounds of 1.29 candy
14-3x=pounds of 1.79 candy
..
2.39x+1.29*2x+1.79(14-3x)=23.86
2.39x+2.58x+25.06-5.37x=23.86
2.39x+2.58x+25.06-5.37x=23.86
.4x=1.2
x=3
2x=6
14-9=5
Pounds to be mixed:
$2.39 candy: 3 pounds
$1.29 candy: 6 pounds
$1.79 candy: 5 pounds